Circular No. 8110

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V838 MONOCEROTIS M. Orio, Istituto Nazionale di Astrofisica, Torino, and University of

Wisconsin; S. G. Starrfield, Arizona State University; and E. Tepedenlegliolu, University of Wisconsin, report that V838 Mon was observed with Chandra (+ ACIS-S) on Feb. 23 for 6800 s, and it was not detected with a 3σ upper limit to the count rate of ≈ 0.04 count/s: “This translates into an upper limit to the unabsorbed flux in the band 0.2–10 keV of F_{x} ≤ 6.5 × 10^{−14} erg cm^{−2} s^{−1}, assuming N_{H} = 4.76 × 10^{21} cm^{−2} [derived from E(B − V) = 0.8 mag; Kimesweger et al. 2002, MNRAS 336, L43] and a plasma emitting at temperature 1 keV, or F_{x} ≤ 4.5 × 10^{−9} erg cm^{−2} s^{−1}, assuming instead a blackbody emitting with a temperature of 30 eV. Given a distance in the range 2000–6000 pc (Bond et al. 2003, Nature, in press), the above upper limit to the x-ray flux rules out that either the central source, or a wind from it, emits x-ray luminosity at the level observed in symbiotic novae a year after the outburst (see Muerset et al. 1997, A.Ap. 319, 201). This is strong evidence against a nova-like thermonuclear runaway as mechanism for the 2002 outburst of this object.”

SUPERNOVA 2003bg IN MCG –05-10-15 D. Pooley and W. H. G. Lewin, Massachusetts Institute of Technology,

on behalf of a larger collaboration, report the detection of x-ray emission from SN 2003bg with the Chandra X-ray Observatory. A 50500-s ACIS-S3 observation on Mar. 24.49 UT reveals a 560-count source (0.5–10 keV) at α = 4^{h}10^{m}59^{s}.42, δ = −31^{o}24′50′′.3 (equinox 2000.0; ± 0′′.5 in α and δ), in close agreement with the optical position (cf. IAUC 8082). Preliminary spectral analysis indicates much low-energy (< 1 keV) emission, perhaps indicating little absorption. A power-law fit gives a photon index of 1.9, but the data are better represented by hot plasma models (MEKAL models in XSPEC) and are equally well fit by a two-component MEKAL model with solar elemental abundances (kT_{1} = 0.5 keV; kT_{2} = 6.5 keV) or a single MEKAL plasma (kT = 5.7 keV) with significant overabundances of O, Ne, and Ni. Both models give an unabsorbed 0.5–10-keV luminosity of ∼ 4 × 10^{39} erg/s (19.1 Mpc, H_{0} = 70 km/s/Mpc). A second Chandra observation is planned in late May. Further optical and radio observations are urged.

GRB 030329 Additional magnitudes from J. McGaha (cf. IAUC 8107): Apr. 5.250

UT, R = 18.39; 5.252, V = 18.34; 5.255, B = 18.83; 5.257, I = 17.95.

2003 April 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green